OpenStudy (anonymous):

What are the solutions to the Quadratic Equation? 2x^2-16x+32=0

hartnn (hartnn):

factor out the '2' first.

OpenStudy (jadeishere):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ }\] a = 2 b = -16 c = 32

OpenStudy (jadeishere):

sorry here's the eqation, forgot to type in "2a"

OpenStudy (jadeishere):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (jadeishere):

if you plug in the values of a, b and c, you should be able to get a fairly simple equation out of that ^.^

hartnn (hartnn):

factor method will be easier than using the formula

OpenStudy (anonymous):

well can you help? im confused. it said to use the formula

OpenStudy (jadeishere):

@hartnn factoring is easier when its an expression, not an equation

OpenStudy (jadeishere):

Yes, I'll help :)

OpenStudy (solomonzelman):

factor out of 22, complete the square !

OpenStudy (solomonzelman):

out of 2 I mean

hartnn (hartnn):

if it explicitly mentions to use formula, then follow what jade says :)

OpenStudy (anonymous):

@SolomonZelman i hate completing the square. lol it makes me confused

OpenStudy (solomonzelman):

dang...

OpenStudy (jadeishere):

That made me feel smart >.<

OpenStudy (mathmale):

Glad that jadeishere is able to help. Note that we could have gotten started out faster had you included all of the instructions (use the quadratic formula).

OpenStudy (solomonzelman):

@ShelbyRenaebb , actually, it is very good, you just have to know when to use it and when not to.

OpenStudy (jadeishere):

Okay, instead of discussing what is easier, can we help? Sorry if that sounds rude :O

OpenStudy (anonymous):

helpppp

OpenStudy (jadeishere):

So in using the quadratic formula, once you plug in the values, it will look like this \[x = \frac{ -(-16)\pm \sqrt{(-16)^2 - 4(2)(32)} }{ 2(2) }\]

OpenStudy (anonymous):

i have five problems i dont know how to do D:

hartnn (hartnn):

shelby, at any point if you don't understand something, do ask doubts

OpenStudy (jadeishere):

A negative multiplied by a negative equals a positive. Meaning -16 will then turn into +16. We'll work on this one first. I'll help with any that I know. :)

OpenStudy (mathmale):

2x^2-16x+32=0 can be done immediately with the quadratic formula, if that's what we're supposed to use to solve it. Then, as someone else has already pointed out, a=a, b=-16 and c=32. Plugging these values into the quadratic formula (given above):\[x=\frac{ -(-16)\pm \sqrt{((-16)^2-4(2)(32))} }{ 2(2) }\] Shelby, can you take the solution friom here?

OpenStudy (anonymous):

Thanks. Just please can we all work together and figure out these five problems? D: I have a test coming up soon and I am studying my study guide and cant figure out these 5 problems

OpenStudy (anonymous):

im doing the quad form now

OpenStudy (jadeishere):

Kay i'll work you through the steps if you need help

hartnn (hartnn):

for quadratic formula, you first need to find values of a,b,c are you comfortable with finding these values ?

OpenStudy (anonymous):

im working them out now

OpenStudy (mathmale):

Hints: -(-16)=16; (-16)^2=256; and (4)(2)(32)=256. Shelby?

OpenStudy (anonymous):

i am at x=16+/- sqroot 16128 over 4

OpenStudy (jadeishere):

\[x = \frac{ 16 \pm \sqrt{256 - 256} }{ 4 }\]

OpenStudy (mathmale):

The 16 and the 4 are fine. Please re-examine what you did to obtain Sqrt(16128). Hint: 256-256=?

OpenStudy (anonymous):

wait id be zero

OpenStudy (jadeishere):

Yep :) did you multiply 256 * 256?

OpenStudy (anonymous):

i thought itd be 0

OpenStudy (mathmale):

Exactly. 256-256=0. And so your TWO roots are real, equal, and equal to what number?

OpenStudy (anonymous):

4?

OpenStudy (jadeishere):

16-0 / 4 = x and 16 + 0 /4 = x

OpenStudy (mathmale):

\[x=\frac{ 16\pm0 }{ 4 }=?\] Yes! your two real, equal roots are {4,4}.

OpenStudy (anonymous):

ya i got 4

OpenStudy (anonymous):

What is the solution to this equation? 5x^2=50

OpenStudy (mathmale):

Just remember that you really do have 2 roots, because you're solving a 2nd-order equation; your 2 roots happen to be equal.

OpenStudy (jadeishere):

yep :) really what the quadratic formula is following PEMDAS, in a certain way

OpenStudy (mathmale):

Would you mind posting this as a new question? (Relenting.) Since you're short on time, let's move forward. Div. both sides by 2. Try to figure out why we wuld do that>

OpenStudy (anonymous):

divide by two??????

hartnn (hartnn):

he meant divide both sides by 5

OpenStudy (anonymous):

lol ok i was like Huhhh?

OpenStudy (anonymous):

x^2=10?

hartnn (hartnn):

yup, now take square root on both sides

OpenStudy (mathmale):

... remembering to use the "plus or minus" sign on the right side of your equation. (Sorry about that "divide by 2" blunder.)

OpenStudy (anonymous):

Haha its okay, im just glad your helping it means alot. I'm trying to get All A's in my classes :D

OpenStudy (anonymous):

so 3.16?

OpenStudy (mathmale):

Glad you're happy! It's nice to feel appreciated.\[\sqrt{x^2}\pm \sqrt{10}\]becomes what?

OpenStudy (anonymous):

x=square root of 10?????

OpenStudy (mathmale):

3.16 is technically correct, but it's much nicer and more accurate to leave your answer in the form\[x=\pm \sqrt{10}\]

OpenStudy (mathmale):

Yes, plus or minus the square root of 10. Again, remember, Shelby: you have TWO roots because your equation is a QUADRATIC (2nd order) one.

OpenStudy (anonymous):

oh ok good to know i should pretty my ending answer lol (:

OpenStudy (mathmale):

Yes, as pretty and ask accurate as possible. We could also color it orange.

OpenStudy (anonymous):

Hahaha

OpenStudy (mathmale):

Shelby: Why not post your next question, but separately? You're doing fine.

OpenStudy (jadeishere):

If I hadn't already given my medal to hartnn i'd give you one @mathmale

OpenStudy (anonymous):

4x^2-6x-10=0 Solve using Quadratic Formula! I think i can do this one but, can we do it together so i can practice???

OpenStudy (anonymous):

Cause its easier on here.

OpenStudy (jadeishere):

Use the equation we had used before Can you tell me which numbers are the a, b and c values?

OpenStudy (anonymous):

i already did the whole problem pretty much, but i got x = 6 +/- sqroot 196 over 8

OpenStudy (jadeishere):

Okay let's start over 4x^2-6x-10=0 if the value of a is 4, what are the values of b and c?

OpenStudy (anonymous):

i got 2.5,-1

OpenStudy (anonymous):

-6 and -10

OpenStudy (jadeishere):

So plugging those into the formula we get *drumroll* \[x = \frac{ -(-6) \pm \sqrt{-6^2 - 4(4)(-10)} }{ 8 }\]

OpenStudy (jadeishere):

\[x = \frac{ 6 \pm \sqrt{36 + 160} }{ 8 }\]

OpenStudy (anonymous):

so i was right

OpenStudy (jadeishere):

Did you have that, at this point?

OpenStudy (jadeishere):

\[x = \frac{ 6 \pm 14 }{ 8 }\]

OpenStudy (anonymous):

yuup!

OpenStudy (jadeishere):

Haha okay x = -1 and x = 2.5 Yes, you're right >.< Nice job :)

OpenStudy (anonymous):

yay!

OpenStudy (jadeishere):

Need any more help?

OpenStudy (anonymous):

Simplify the imaginary unit i. Square root of -16

OpenStudy (jadeishere):

okay that requires complex numbers, not my strongest area. @AravindG @beccaboo333

OpenStudy (anonymous):

ok ill open it in neww question

OpenStudy (jadeishere):

Kaykay